### v3

# Variational Learning Finds Flatter Solutions at the Edge of Stability

Abstract
Variational Learning (VL) has recently gained popularity for training deep neural networks. Part of its empirical success can be explained by theories such as PAC-Bayes bounds, minimum description length and marginal likelihood, but little has been done to unravel the implicit regularization in play. Here, we analyze the implicit regularization of VL through the Edge of Stability (EoS) framework. EoS has previously been used to show that gradient descent can find flat solutions and we extend this result to show that VL can find even flatter solutions. This result is obtained by controlling the shape of the variational posterior as well as the number of posterior samples used during training. The derivation follows in a similar fashion as in the standard EoS literature for deep learning, by first deriving a result for a quadratic problem and then extending it to deep neural networks. We empirically validate these findings on a wide variety of large networks, such as ResNet and ViT, to find that the theoretical results closely match the empirical ones. Ours is the first work to analyze the EoS dynamics of VL.
1 Introduction
Variational Learning (VL) has been used to perform deep learning from early on (Graves, 2011; Blundell et al., 2015) and recently also started to show good results at large scale. It has been shown to outperform state-of-the-art optimizers without any increase in the cost. For example, on ImageNet, VL substantially improves overfitting commonly seen in AdamW and for pretraining GPT-2 from scratch, VL achieves a lower validation perplexity than AdamW (Shen et al., 2024). For low-rank fine-tuning of Llama-2 (7B), VL improves both accuracy (by $2.8\%$) and calibration (by $4.6\%$) (Cong et al., 2024; Li et al., 2025). Moreover, VL methods explicitly derived by using PAC-Bayes bounds (Wang et al., 2023b; Zhang et al., 2024b) have shown consistent improvements over AdamW. All such results confirm the importance of VL for deep learning.
Despite these successes, the theoretical mechanisms behind the good performance of VL remain poorly understood. It is often assumed that simplistic Gaussian posteriors, such as those used currently for deep learning, may not be enough because they are poor approximations of the true posterior; lack of a good prior is another issue. Despite these concerns, VL shows good performance in practice. Theories such as minimum description length (Hinton and Van Camp, 1993; Hochreiter and Schmidhuber, 1997; Blier and Ollivier, 2018), PAC-Bayes (Dziugaite and Roy, 2017; Zhou et al., 2019; Lotfi et al., 2024; Alquier et al., 2024) and marginal likelihood (Smith and Le, 2017; Immer et al., 2021) can partially explain the success. In particular, PAC-Bayes theory provides a natural explanation for why flatter solutions may generalize better, see for instance the discussion in (Alquier et al., 2024, Section 3.3). However these theories say little about the regularizing properties of the learning algorithm. Similarly to deep learning, the presence of implicit regularization is likely also at play in VL, but few tools exist to unravel these effects.
In this work, we analyze the implicit regularization of VL algorithms at the Edge of Stability. The EoS analysis has previously been used to show that Gradient Descent (GD) with constant learning rate $\rho$ implicitly biases the trajectories towards flatter solutions, where the sharpness (defined as the operator norm of the loss Hessian) hovers around $2/\rho$. We extend this analysis to VL and show that sharpness can be further lowered by controlling the posterior covariance and the number of Monte-Carlo samples used to compute posterior expectations; see an illustration in Figure 1(a). Similarly to the standard EoS technique, we first derive an exact expression for the stability threshold for VL on a quadratic problem and then propose extensions to general loss functions. We then empirically validate these finding on a wide variety of deep networks, including Multi-Layer Perceptron, ResNet, and Vision Transformers; see an example in Figure 1(b). We observe similar results when posterior shape is automatically learned, for instance, by using diagonal covariance Gaussians and heavy-tailed posteriors. Code to replicate these results is available at https://github.com/Avra98/variationallearning-eos.
Figure 1: Panel (a): The left figure shows trajectory traces of VL on a quadratic problem with an isotropic variational posterior whose mean is learned but variance is set to a fixed value. The trajectory becomes more unstable as the posterior variance is increased and number of Monte-Carlo samples is decreased. We provide an exact expression to compute the stability threshold at which the iterations become unstable (Theorem 5). Panel (b): We show the validity of the threshold on neural network training. The right figure (top) shows this on CIFAR-10 for an MLP where VL achieves lower sharpness than GD when posterior variance is increased. The bottom figure shows that the sharpness (solid line) matches the stability threshold obtained by our theorem (dashed line).
2 Theoretical Tools to Analyze Variational Learning
Variational Learning optimizes the variational reformulation of Bayesian learning (Zellner, 1988), where the goal is to find good approximations of the Gibbs distribution $\exp(-\ell(\boldsymbol{\theta}))/\mathcal{Z}$ with partition function $\mathcal{Z}$ over parameters $\boldsymbol{\theta}\in\mathbb{R}^{d}$. Specifically, we seek the closest distribution $q(\boldsymbol{\theta})$ in a set of distributions $\mathcal{Q}$ that minimizes the expected loss regularized by the entropy:
$\mathop{\mathrm{arg\,min}}_{q\in\mathcal{Q}}\;\mathbb{E}_{\boldsymbol{\theta}\sim q}[\ell(\boldsymbol{\theta})]-\mathcal{H}(q).$ (1)
This is an instance of the maximum-entropy principle. The objective is equivalent to minimizing the KL divergence to the Gibbs posterior and naturally encourages higher-entropy (flatter) solutions due to the entropy $\mathcal{H}(q)$; for example, see the illustrative example in Khan and Rue (2023, Figure 1). Similar arguments can also be made with the minimum-description-length principles (Hinton and Van Camp, 1993; Graves, 2011; Blier and Ollivier, 2018). In practice, Variational Learning has started to show good results achieving generalization performance better than the state of the art optimizers across several tasks (Shen et al., 2024; Cong et al., 2024). This matches with the intuition: Because the variational objective prefers wider distributions, we expect the posterior to be located in the region where the loss $\ell(\boldsymbol{\theta})$ is flatter, see Figure 2.
Despite this intuition, there is little work to analyze the implicit regularization that could help us understand how VL favors flatter regions and how we can control it. Existing theories do not sufficiently address this; a review of the related work is in Appendix A. We suspect the implicit regularization to be related to the shape of the posterior but currently there are no results explicitly characterizing this.
Figure 2: VL’s mechanism for flatter minima: The posterior variance determines the minima’s location. A small variance settles the posterior in a sharp minima (left), while a larger variance allows it to explore and find a flat minima (right).
VL is also closely connected to weight-noise or weight-perturbation methods where the goal is to optimize $\mathbb{E}_{q(\boldsymbol{\epsilon)}}[\ell(\boldsymbol{\theta}+\boldsymbol{\epsilon)}]$ with $q(\boldsymbol{\epsilon)}$ being a fixed distribution to inject weight noise. This can be seen as a special case of variational learning where the shape of the posterior is fixed and only the location is learned. Multiple works (Zhu et al., 2019; Nguyen et al., 2019; Zhang et al., 2019; Jin et al., 2017; Simsekli et al., 2019) have analyzed generalization behavior of such weight-noise variational methods but, even for this simple case, there are no studies connecting them to the EoS results for GD. In this paper, we will address these gaps and provide both theoretical and empirical results regarding the edge of stability phenomenon of such a variational GD (VGD).
2.1 Edge of Stability for Gradient Descent
We will briefly review the EoS result for GD. The standard EoS literature relies on a result for a quadratic problem and then extends it to deep neural networks. For instance, consider the following
$\text{quadratic loss:}\quad\ell(\boldsymbol{\theta})=\frac{1}{2}\boldsymbol{\theta}^{\top}\mathbf{Q}\boldsymbol{\theta},\quad\text{where}\quad\mathbf{Q}=\sum_{i=1}^{d}\lambda_{i}\mathbf{v}_{i}\mathbf{v}^{T}_{i}.$ (2)
Here, $\mathbf{Q}$ is a positive definite matrix with $\lambda_{i}$ being its $i^{th}$ largest eigenvalue and $\mathbf{v}_{i}$ being the corresponding eigenvector. The following result states the condition under which one step of GD leads to a decrease in the loss.
Lemma 2.1.
(Descent Lemma) For a GD update $\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}-\rho\nabla\ell(\boldsymbol{\theta}_{t})$ on the quadratic loss (2), the loss decreases at each step, that is, we have
$\ell(\boldsymbol{\theta}_{t+1})-\ell(\boldsymbol{\theta}_{t})\leq 0,\quad\text{if and only if}\quad\lambda_{i}\leq\frac{2}{\rho}\quad\text{for all }i.$ (3)
This lemma implies that GD converges on a quadratic loss if all eigenvalues satisfy $\lambda_{i}<2/\rho$. This is a different way of writing the standard condition that maximum eigenvalue $\lambda_{1}<2/\rho$. The result extends to any smooth $\ell(\boldsymbol{\theta})$ and for such cases, the condition implies that the learning rate $\rho<2/\beta$, where $\beta=\sup_{\boldsymbol{\theta}}\|\nabla^{2}\ell(\boldsymbol{\theta})\|_{2}$ is the bound on the Hessian norm (Nesterov et al., 2018, Chapter 2). This condition is necessary and sufficient for descent of GD on quadratic.
While the descent lemma is predictive of convergence of GD for smooth functions, deep learning exhibits a more complex behavior. When training deep neural networks with a constant learning rate $\rho$, the Hessian’s operator norm (or sharpness) tends to settle around the value $2/\rho$. This phenomenon is referred to as the Edge of Stability. Deep neural networks often operate at this edge and converge in a non-monotonic, unstable manner. A comprehensive empirical investigation of this phenomenon is given by Cohen et al. (2021), where two phases of training neural networks were noticed. In the first phase referred to as ‘progressive sharpening’, the Hessian’s operator norm, $\|\nabla^{2}\ell(\boldsymbol{\theta}_{t})\|_{2}$ increases and slowly approaches $2/\rho$. This is followed by the second, EoS phase, where sharpness hovers around $2/\rho$ and the loss continues to decrease in an oscillatory fashion.
This phenomenon does not happen on a quadratic loss, but only for certain losses with nonzero third-order derivative. As shown by Damian et al. (2023), once sharpness reaches $2/\rho$, a local quadratic approximation is insufficient to capture the dynamics because the third order Taylor expansion term of the loss becomes significant. This cubic term represents the gradient of the sharpness, which serves as a negative feedback to counteract progressive sharpening and stabilize the sharpness around $2/\rho$ in GD. Instead of divergence, the iterates exhibit oscillatory or non-monotonic behavior even when the sharpness reaches $2/\rho$. This is the reason why $2/\rho$ is also called the ‘Stability Threshold’. Increasing $\rho$ reduces the edge, which could then drive the iterates toward flatter minima that may generalize better. EoS analysis can help us understand such implicit regularization during training, especially for nonconvex problems.
Different optimizers have different stability thresholds which depends on the sharpness value $\|\nabla^{2}\ell(\boldsymbol{\theta}_{t})\|_{2}$. For instance, Sharpness Aware Minimization (SAM) leads to a different, smaller stability threshold (Long and Bartlett, 2024) compared to $2/\rho$ for GD. In this work, we show a similar result where VL has a smaller threshold than GD. We show this by deriving the stability threshold for a simpler quadratic problem first and analyze several factors such as posterior covariance and the number of posterior samples that influence the threshold. Then, we empirically demonstrate that similar results hold for the case when VL is used to train deep neural networks.
3 Stability Threshold for a Simple Case of Variational Learning
We start with a simple VL setting where the goal is to estimate a Gaussian $q(\boldsymbol{\theta})=\mathcal{N}(\boldsymbol{\theta}|\,\boldsymbol{m},\boldsymbol{\Sigma})$ whose mean $\boldsymbol{m}$ is unknown and covariance $\boldsymbol{\Sigma}$ is fixed. We assume $\boldsymbol{\Sigma}=\sigma^{2}\mbox{$\mbox{$\mathbf{I}$}$}$ is an isotropic covariance matrix, with scalar variance $\sigma^{2}$. Khan and Rue (2023) [Section 1.3.1] show that this can be estimated by the following Variational GD algorithm:
$\mathbf{m}_{t+1}\leftarrow\mathbf{m}_{t}-\rho\,\mathbb{E}_{\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})}[\nabla\ell(\mathbf{m}_{t}+\boldsymbol{\epsilon})].$ (4)
This algorithm can be implemented by the following version where the expectation is estimated by drawing $N_{s}$ Monte Carlo samples to approximate the expectation as follows
$\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1}\leftarrow\mathbf{m}_{t}-\rho\,\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\nabla\ell(\mathbf{m}_{t}+\boldsymbol{\epsilon_{i}}),\quad\boldsymbol{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mbox{$\mbox{$\mathbf{I}$}$}).$ (5)
The updated iterate $\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1}$ depends on the $N_{s}$ Monte Carlo samples $\boldsymbol{\epsilon}=[\boldsymbol{\epsilon}_{1},\boldsymbol{\epsilon}_{2},..,\boldsymbol{\epsilon}_{N_{s}}]$. Compared to the standard gradient descent, the update step in Variational GD (VGD), is determined by gradients averaged over a local neighborhood of perturbed weights. As a result, the update introduces two interacting effects influencing its stability threshold: (1) a perturbation effect, originating from the perturbation covariance $\boldsymbol{\Sigma}$, and (2) a smoothing effect, resulting from averaging gradients across $N_{s}$ Monte Carlo samples. Similarly to Lemma 2.1, we can derive the stability threshold for the above VGD and show that it is smaller than that of GD. {restatable}theoremmainproof Consider the VGD update (5) on the quadratic loss (2), then
$\mathbb{E}_{\boldsymbol{\epsilon}}[\ell(\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1})]-\ell(\mathbf{m}_{t})<0\quad\text{if}\quad\lambda_{i}<\frac{2}{\rho}\cdot\operatorname{VF}\left(\frac{N_{s}}{\sigma^{2}}\cdot c_{i,t}\right)\quad\text{for all }i,$ (6)
where $c_{i,t}=(\lambda_{i}\mathbf{m}_{t}^{\top}\mathbf{v}_{i})^{2}$, and $\operatorname{VF}(\cdot)$ denotes the Variational Factor given by
$\operatorname{VF}(z):=\rho\cdot\sqrt{\frac{z}{3}}\cdot\sinh\left(\frac{1}{3}\operatorname{arcsinh}\left(\frac{3}{\rho}\sqrt{\frac{3}{z}}\right)\right).$ (7)
See Appendix B for a proof where we also discuss the case where $\mathbf{Q}$ is low rank. The theorem is analogous to the descent lemma for GD and it states that the Variational GD (5) also decreases the expected loss over $\boldsymbol{\epsilon}$ if each eigenvalue $\lambda_{i}$ is less than $2/\rho$ times a function called the Variational Factor (VF). The VF function is strictly less than $1$, therefore the stability threshold is strictly less than $2/\rho$. Unlike GD, the condition here is only sufficient but not necessary. A necessary condition can also be derived but the one above is sufficient for this paper.
Figure 3(a) shows the stability threshold as a function of $z$ where it is clear that it always lower bounds $2/\rho$ (dashed horizontal line). The exact value of VF depends on its argument $z$ which in Equation˜5 mainly depends on the number of Monte-Carlo samples $N_{s}$ and posterior variance $\sigma^{2}$, but also on the loss-dependent constant $c_{it}$ which is essentially a function of $\lambda_{i}$ and $\mathbf{m}_{t}^{\top}\mathbf{v}_{i}$.
Theorem 1 guarantees a decrease in the expected loss. Below, we state another result to show that the actual loss also decreases with high probability if the expected loss decreases by a margin $\delta>0$. {restatable}lemmadescentconc In the same setting as Equation˜5, when the expected loss at next iteration is smaller than the previous loss by some margin $\delta>0$, that is,
$\mathbb{E}_{\boldsymbol{\epsilon}}[\ell(\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1})]<\ell(\mathbf{m}_{t})-\delta,$
then $\ell(\mathbf{m}^{\boldsymbol{\epsilon}}_{t+1})-\ell(\mathbf{m}_{t})<0$ occurs with probability at least $1-2\exp\left(-c_{1}\min\left\{\delta^{2}N_{s}^{2}/c_{2},\delta N_{s}/c_{2}\right\}\right)$, for constants $c_{1},c_{2}>0$ depending only on $\rho$, $\mathbf{Q}$, and $\boldsymbol{\Sigma}$. The result also shows that the probability with which this happens increases with the number of Monte Carlo samples $N_{s}$.
(a) Variational Factor (VF)
(b) Descent probability
(c) Stability
Figure 3: (a) Solid black curve shows the theoretical stability-threshold of VGD as a function of $N_{s}/\sigma^{2}$. The curve is clearly lower than the stability threshold of GD, shown with the horizontal dashed, gray line. (b) Empirical verification on a scalar quadratic problem with curvature $\lambda$ where we plot the empirically computed probability of descent for VGD runs with different values of $\sigma^{2}$. We show a heatmap for $(\lambda,1/\sigma^{2})$ values where lighter colors indicate higher probability of descent. We overlay the heatmap with the stability thresholds of VGD (red solid curve), clearly showing that theoretical limit shown in (6) matches the empirical probability. (c) The figure further includes $N_{s}$ and marks the region where a pair $(N_{s},\sigma^{2})$ will either lead to descent or not (marked with ‘stable’ and ‘unstable’ respectively).
So far, we show that the stability threshold for Variational GD is smaller than that of GD, but can we also ensure it to be smaller in practice? The answer is yes, and it can be done by controlling $\sigma^{2}$ and $N_{s}$. We will now demonstrate this on a 1D quadratic loss $\ell(m)=\mbox{$\frac{1}{2}$}\lambda m^{2}$. For such cases, GD with step-size $\rho$ descends whenever the curvature $\lambda$ is bounded by $2/\rho$, but we can show that VGD descends only for lower curvature values. To show this, we run VGD for many $(\lambda,\sigma)$ pairs. For each run, we use 10 random realizations of $\epsilon$, then record how often the loss decreases, and finally compute the approximate descent probability $\mathbb{P}\big(\ell(m_{t+1}^{\epsilon})<\ell(m_{t})\big)$.
Figure 3(b) shows this probability as a heatmap where lighter color indicate a higher probability of descent; a white pixel indicates a probability of 1 and a black one indicates a probability of 0. We overlay this with the solid red curve showing the theoretical stability-threshold of Variational GD as dictated by (6), that is, $\lambda=(2/\rho)\,\mathrm{VF}(N_{s}/\sigma^{2})$. The theoretical curve closely matches the transition boundary where probabilities transition from 1 to 0. The figure clearly shows that by increasing $\sigma^{2}$ we can reduce the stability threshold in practice too. Figure 3(c) further illustrates the effect of changing $N_{s}$, showing the regions where a $(N_{s},\sigma^{2})$ pair lead to descent (marked as ‘stable’) or otherwise (marked with ‘unstable’). As expected, the relationship is linear and the same effect is obtained by either increasing $N_{s}$ or decreasing $\sigma^{2}$.
3.1 Nature of Stability in GD and VGD
The introduction of noise in VGD fundamentally changes the nature of its stability compared to the deterministic behavior of GD. The descent lemma for GD on a quadratic guarantees that the iterates are asymptotically stable, which is defined as follows:
Definition 3.1 (Asymptotic Stability Lyapunov (1992), Chapter 2).
Let $\boldsymbol{\theta}^{*}$ be the minimum. An iterate $\boldsymbol{\theta}_{k}$ is asymptotic stable if it is Lyapunov stable and there also exists a $\delta>0$ such that if the algorithm starts within a $\delta$-neighborhood of the fixed point, the iterate will converge to the fixed point. Formally, if $\|\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}\|<\delta$ for a finite $k$, then:
$\lim_{k\to\infty}\|\boldsymbol{\theta}_{k}-\boldsymbol{\theta}^{*}\|=0.$ (8)
GD on a quadratic is asymptotically stable because the condition learning-rate $\rho<2/\lambda_{1}$ ensures that the iteration matrix $(\mathbf{I}-\rho\mathbf{Q})$ has all eigenvalues less than $1$, guaranteeing convergence of the iterates to its fixed point. Analogously VGD, the iterates are Stochastically Stable which is defined as follows
Definition 3.2 (Stochastic Stability Kushner (1972), Chapter 2).
Let $\boldsymbol{\theta}^{*}$ be the minimum. The VGD iterates $\boldsymbol{\theta}^{\boldsymbol{\epsilon}}_{t}$ is said to be stochastically stable in the mean-square sense if there exists a constant $C>0$ such that for any initial point $\boldsymbol{\theta}_{0}$, the iterates $\boldsymbol{\theta}^{\boldsymbol{\epsilon}}_{t}$ satisfy:
$\mathbb{E}_{\boldsymbol{\epsilon}}\left[\|\boldsymbol{\theta}^{\boldsymbol{\epsilon}}_{t}-\boldsymbol{\theta}^{*}\|^{2}\right]\leq C\|\boldsymbol{\theta}_{0}-\boldsymbol{\theta}^{*}\|^{2},\quad\text{for all }t>0.$ (9)
This form of stability ensures that the iterates remain bounded in the mean-square error sense, preventing them from diverging. Practically, this corresponds to the behavior where the distribution of the iterates converges to a stationary distribution, rather than the iterates themselves converging to a single fixed point as in asymptotic stability. The condition for probabilistic descent, established in Theorem 5, is what characterizes this stable behavior. By ensuring the loss decreases on average, it prevents the divergence of the iterates, confining them to a stable random walk that converges in distribution to a stationary state around the minimum. Accordingly, any reference to "stability" or a "stability threshold" throughout for VGD in this paper refers to this notion of stochastic stability.
3.2 Comparison with Regularization Effect of SGD
Our work differs fundamentally from studies on the regularization effects of mini-batch noise in SGD, such as Wu et al. (2022); Zhu et al. (2019); Wu et al. (2018); Ibayashi and Imaizumi (2023); Mulayoff and Michaeli (2024), which analyze gradient noise and its role in promoting flatter solutions. In contrast, VGD introduces noise directly in the weights, inducing a structured and anisotropic form of gradient noise shaped by the curvature of the loss landscape, something not captured by existing SGD analyses. While most SGD-based studies assume Gaussian gradient noise, despite empirical evidence of heavy-tailed behavior (Gurbuzbalaban et al., 2021; Nguyen et al., 2019), we show that for the quadratic loss, Gaussian perturbations in weights lead to Gaussian-distributed gradients, that is, $\hat{\mathbf{g}}\sim\mathcal{N}(\nabla\ell(\mathbf{m}_{t}),\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q})$, where the covariance is amplified along sharper directions. This formulation requires no assumptions on the shape of gradient noise (unlike, e.g., Lee and Jang (2023)). To complement our work, we further perform an empirical study using weight perturbations drawn from a heavy-tailed distribution in deep neural networks.
(a) ViT
(b) MLP
(c) ResNet-20
Figure 4: Smaller sharpness corresponds to higher test accuracy for network architectures trained on CIFAR-10. Panels show (a) ViT, (b) MLP, and (c) ResNet-20. Full batch GD with Variational GD achieves lower sharpness and better test accuracy.
4 Experiments on Deep Neural Networks
Similarly to the GD case, we expect the stability threshold for the quadratic to serve as an EoS limit. That is, we expect that, when using variational learning for deep neural networks, the sharpness should hover around the stability limit. By controlling the covariance shape and number of Monte Carlo samples we can stere optimization into regions where sharpness is much lower than GD. The following hypothesis formally states this intuition which we verify through extensive experiments.
Hypothesis 1.
For a twice-differentiable loss $\ell(\mathbf{m})$ optimized using the update rule (5), the top Hessian eigenvalue $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$ hovers around the stability threshold,
$\frac{2}{\rho}\cdot\mathrm{VF}\left(\frac{N_{s}}{\sigma^{2}}\cdot c_{1,t}\right),$
where $c_{1,t}=(\mathbf{v}_{1,t}^{\top}\nabla\ell(\mathbf{m}_{t}))^{2}$ and $\mathbf{v}_{1,t}$ denotes the top eigenvector of the Hessian $\nabla^{2}\ell(\mathbf{m}_{t})$, and $\nabla\ell(\mathbf{m}_{t})$ is the gradient evaluated at $\mathbf{m}_{t}$.
To validate the hypothesis, we conduct extensive experiments on standard architectures (MLPs, ResNets) and modern ones such as Vision Transformers. In Figure 4, we plot the full training dynamics of Variational GD, including test accuracy, training accuracy, and sharpness. Across all three architectures on the CIFAR-10 classification task using MSE loss, Variational GD with isotropic Gaussian noise consistently achieves lower sharpness and higher test accuracy compared to GD.
(a) $\rho=0.01$
(b) $\rho=0.02$
(c) $\rho=0.05$
Figure 5: Normalized Sharpness $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}/(2/\rho)$ hovers around the Variational Factor in MLP.
In Figure 5, we plot the normalized sharpness ($\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$ divided by $2/\rho$) for MLPs, evaluated at each training step, alongside the corresponding Variational Factor (VF) for various learning rates $\rho$ and posterior variances $\sigma^{2}$. The results support Hypothesis 1, showing that the normalized sharpness closely tracks the predicted VF across settings. This further validates the use of a local quadratic approximation to analyze stability dynamics, consistent with several prior works. Unlike the stability analysis for GD on a quadratic where the threshold $2/\rho$ is constant, the stability condition in VGD is dynamic since it depends on the gradient evaluated at the mean. This explains the fluctuation of the stability threshold across iterations in VGD. In the Appendix (Figure 14), we demonstrate that a similar phenomenon holds for ResNet-20. Next, we isolate and examine the roles of the posterior variance and the number of samples in determining this stability threshold.
Perturbation effect of Gaussian variance: In Figure 6, we present experiments on VGD for a classification task using a multi-layer perceptron (MLP), where we plot the sharpness of the final iterate as a function of the variance $\sigma^{2}$ of an isotropic Gaussian. For three different learning rates, $\rho=0.05$, $0.1$, and $0.2$, we run VGD with different $\sigma^{2}$. Across all learning rates $\rho$ and numbers of posterior samples $N_{s}$, we observe that larger variance consistently leads to lower sharpness. This exactly aligns with Hypothesis 1, larger perturbations reduce the stability threshold, promoting escape from sharper regions of the loss landscape.
Smoothing effect of posterior samples $N_{s}$: Larger posterior sample size $N_{s}$ reduces the variance of the perturbed gradient estimator $\hat{\mathbf{g}}=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\nabla\ell(\mathbf{m}_{t}+\boldsymbol{\epsilon}_{i})$, which satisfies $\hat{\mathbf{g}}\sim\mathcal{N}(\nabla\ell(\mathbf{m}_{t}),\smash{\frac{1}{N_{s}}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q})$ under a quadratic loss. As $N_{s}$ decreases, the variance increases (scaling as $1/N_{s}$), lowering the stability threshold and enabling escape from sharper regions. This effect is shown in Figure 6, where smaller $N_{s}$ consistently yields lower sharpness across learning rates and variances. Additional discussion on the role of posterior samples, is provided in Appendix C.
(a) $2/\rho=40$
(b) $2/\rho=20$
(c) $2/\rho=10$
Figure 6: Sharpness of the final iterate for an MLP trained on CIFAR-10 for variational GD across learning rate $\rho$, noise covariance $\boldsymbol{\Sigma}$, and posterior samples $N_{s}$. Higher variance and smaller samples lead to lower sharpness. For panel (c), training did not converge for variance values 2 and 5, therefore not shown in the plot.
4.1 Edge of Stability under Heavy-Tailed Noise
In variational gradient descent (VGD), the stability threshold depends not only on the perturbation covariance and sample size, but also on the posterior’s tail behavior. We empirically investigate this by drawing weight perturbations from a Student-t distribution with degrees of freedom $\alpha$, where smaller $\alpha$ induces heavier tails and larger deviations in the perturbed gradients. The distribution becomes heavier-tailed as $\alpha$ decreases, with the Gaussian recovered as $\alpha\to\infty$. In Figure 7, we train an MLP on CIFAR-10 and observe that heavier-tailed perturbations yield lower sharpness and better test accuracy. Similar trends are confirmed for ResNet-20 and ViT in Appendix F. While a theoretical analysis under heavy-tailed noise is challenging, our results highlight that the posterior shape plays a critical role in generalization.
Figure 7: Sharpness dynamics with noise injection from a heavy tailed Student t posterior parameterized by $\alpha$. Perturbations from heavier tails (smaller $\alpha$) lead to smaller sharpness.
4.2 Adaptive Edge of Stability and Variational Online Newton Methods
Recent work by Cohen et al. (2022) studies the dynamics of adaptive gradient methods, introducing the concept of Adaptive Edge of Stability (AEoS). In this section, we will show that a similar phenomenon happens in natural-gradient variational learning.
Cohen et al. (2022) show that for adaptive gradient methods instead of sharpness $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$, a modified quantity, termed preconditioned sharpness $\|\text{diag}(\mathbf{p}_{t})^{-1}\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}$ hovers around $2/\rho$, where $\mathbf{p}_{t}$ is a preconditoner. Adaptive gradient methods differ from standard gradient descent as the former can adapt their preconditioner $\mathbf{P}_{t}$ and move into high curvature regions. For example in Adam, a preconditioner $\mathbf{p}_{t}$ is updated in an exponential moving average (EMA) fashion with parameter $\beta_{2}$:
$\mathbf{v}_{t+1}=\beta_{2}\mathbf{v}_{t}+(1-\beta_{2})\nabla\ell(\mathbf{m}_{t})^{2},\;\mathbf{p}_{t+1}=\sqrt{\frac{\mathbf{v}_{t+1}}{1-\beta_{2}^{t+1}}},\;\mathbf{m}_{t+1}=\mathbf{m}_{t}-\rho\,\nabla\ell(\mathbf{m}_{t})/\mathbf{p}_{t+1}.$ (10)
Here, all operations such as squaring or division of vectors are performed elementwise.
Natural-gradient VL methods which learn a complete Gaussian posterior $q_{t}=\mathcal{N}(\mathbf{m}_{t},\mathbf{P}_{t}^{-1})$ take a similar form to the above adaptive gradient methods. An instance of this is the Variational Online Newton (VON) update rule (Khan and Rue, 2023, Eq. 12),
$\mathbf{m}_{t+1}\leftarrow\mathbf{m}_{t}-\rho\,\mathbf{P}_{t+1}^{-1}\,\mathbb{E}_{q_{t}}[\nabla_{\theta}\ell(\boldsymbol{\theta})]\quad\text{and}\quad\mathbf{P}_{t+1}\leftarrow(1-\beta_{2})\,\mathbf{P}_{t}+\beta_{2}\,\mathbb{E}_{q_{t}}[\nabla^{2}_{\theta}\ell(\boldsymbol{\theta})].$ (11)
Here, the posterior covariance $\mathbf{P}_{t}^{-1}$ is learned using the loss Hessians. The variational GD in Eq. (5) corresponds to the special case where $\mathbf{P}_{t}$ is fixed across iterations. Adaptive optimizers such as Adam, RMSProp, and Adadelta can be seen as special cases of VON, as shown in (Khan and Rue, 2023, Section 4.2).
Figure 8: Preconditioned sharpness for Adam and IVON across temperatures $\tau$. Smaller $\tau$ shrinks the posterior and yields larger preconditioned sharpness.
Here, we study the AEoS for a recent large-scale implementation of Equation˜11 by Shen et al. (2024) called IVON (Improved Variational Online Newton). There, the update for the preconditioner is approximated using Stein’s identity to estimate the Hessian from $N_{s}$ samples:
$\mathbb{E}_{q_{t}}[\mbox{$\mbox{diag}$}(\nabla^{2}_{\theta}\ell(\theta))]\approx\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(\nabla\ell(\boldsymbol{\theta}_{i})\cdot\frac{\boldsymbol{\theta}_{i}-\mathbf{m}_{t}}{\sigma^{2}}\right),\quad\boldsymbol{\theta}_{i}\sim q_{t}.$ (12)
In Figure 8, we compare the preconditioned sharpness of IVON and Adam on MLPs by varying the temperature $\tau$, which linearly scales covariance of the posterior distribution $q_{t}$, see (Shen et al., 2024, Algorithm 1, line 8). We observe that IVON consistently yields lower preconditioned sharpness than the $2/\rho$ threshold typically observed in Adam. We conjecture the lower sharpness is due to the noise injection and smoothing effects in estimating both the gradient and curvature. Additional comparisons for ResNet-20 and ViT are provided in Appendix F. Computing the exact stability threshold for IVON is nontrivial, as it requires a joint analysis of the coupled dynamics in Eq. 11, and an interesting direction for future work.
4.3 Effect of Batch Size
For nonconvex problems such as deep learning, minimizing the variational objective (1) is not, by itself, sufficient to guarantee flat minima or a low objective value. The choice of optimization dynamics and hyperparameters play a crucial role as well. We demonstrate this by running IVON with mini-batching across varying learning rates $\rho$ and batch sizes $B$. As shown in Figure 9, larger learning rates and smaller batch sizes consistently lead to better local minima of the variational objective. The present work offers an explanation why large learning rates work better, but we also speculate that smaller batch sizes encourages broader posteriors in flatter minima allowing a smaller objective value. These results further highlight the importance of choice of hyperparameter and optimizers in variational learning. Poor performance of variational methods as claimed in the literature may not be due to flaws in the variational formulation itself but due to unfavorable optimization dynamics or choices of hyperparameters.
(a) Varying learning rate
(b) Varying batch size
Figure 9: Minimizing the variational objective depends on the choice of hyperparameter and its implicit regularization effect, visible in both the objective value and sharpness reduction.
5 Conclusion and Discussion
In this work, we study the regularization effect in Variational Learning (VL) that enables it to find solutions with better generalization than Gradient Descent (GD). We show that the sharpness dynamics in VL can be accurately tracked through its instability mechanism under a local quadratic approximation of the loss. We argue that to fully explain generalization in VL, one must look beyond theoretical frameworks like PAC-Bayes bounds and instead understand the optimization dynamics and the implicit regularization they induce. We hope our work takes a positive step in this direction and motivates further investigation into the role of training dynamics in Bayesian deep learning.
Acknowledgements
This work is supported by the Bayes duality project, JST CREST Grant Number JPMJCR2112. A.Ghosh and R.Wang acknowledge support from NSF Grant CCF-2212065. S.Ravishankar acknowledge support from NSF CAREER CCF-2442240 and NSF Grant CCF-2212065. M.Tao is grateful for partial supports by NSF Grant DMS-1847802, Cullen-Peck Scholarship, and Emory-GT AI.Humanity Award. We acknowledge Zhedong Liu for his help with the experiments on heavy-tailed distribution and Pierre Alquier for helpful discussions.
Impact Statement
This work improves understanding of how learning algorithms can find solutions that generalize better. It may lead to more reliable and robust AI systems. While primarily theoretical, the insights could support better model training in practical settings. We see minimal risk of misuse, but encourage responsible use in sensitive applications.
Supplementary Material
The supplementary material contains the following appendix section.
  • Related Works ........................................................................................................................................................................A
  • Proof of Theorem 1 ........................................................................................................................................................................B
  • Role of posterior samples on a quadratic ........................................................................................................................................................................C
  • Loss Smoothing Beyond Local Quadratic Approximation ........................................................................................................................................................................D
  • Discussion on stability and descent for Variational GD ........................................................................................................................................................................E
  • Additional experiments across diverse datasets ........................................................................................................................................................................F
Appendix A Related Works
Variational Learning
Given a prior and a likelihood, variational inference approximates the posterior by optimizing the evidence lower bound within a family of posterior distribution. The ELBO is optimized by natural gradient descent [Amari, 1998], in the hope that steepest descent induced by the KL divergence is a better metric to compare probability distributions, this gives to rise of a class of algorithms called Natural Gradient Variational Learning. Several works have attempted to apply this to deep learning [Khan et al., 2018, Osawa et al., 2019, Lin et al., 2020]. Recently, Shen et al. [2024] provide an improved version of varlational Online Newton that largely scales and obtains state of the art accuracy and uncertainty at identical cost as Adam. A step of VL includes weight perturbation based on noise injection from the posterior distribution which has similarities to noise injection.
Regularization Effect of Noise
Injecting noise into the parameters within gradient descent has several desirable features such as escaping saddle points [Jin et al., 2017, Reddi et al., 2018] and local minima Zhu et al. [2019], Nguyen et al. [2019]. In fact it has been widely observed empirically that the SGD noise has an important role to play to find flat minima. Noise with larger scale [Zhu et al., 2019, Nguyen et al., 2019, Zhang et al., 2019, Smith et al., 2020, Wei and Schwab, 2019] and heavy-tail [Simsekli et al., 2019, Panigrahi et al., 2019, Nguyen et al., 2019, Wang et al., 2022a] drives the optimization trajectories towards flatter minima. Orvieto et al. [2023] demonstrated through stochastic Taylor expansion that injecting Gaussian noise into parameters before a gradient step implicitly regularizes by penalizing the curvature of the loss, while Orvieto et al. [2022] showed that anticorrelated noise in Perturbed Gradient Descent (PGD) specifically penalizes the trace of the Hessian. Zhang et al. [2024a] showed that injecting noise in opposite directions to the weight space leads to a better regularization of the trace of the Hessian.
Gradient Descent at Edge of Stability
Cohen et al. [2021], building on the work of Jastrzebski et al. [2020] empirically showed that for full batch Gradient Descent (GD) with constant step-size ($\rho$), the operator norm of the Hessian (also termed as sharpness) settles in a neighborhood of $2/\rho$. This threshold is termed as edge of stability (EoS) because gradient descent on a quadratic only converges if the sharpness is below $2/\rho$. Strikingly in complex neural network landscapes, sharpness settling around the stability limit $2/\rho$ (instead of diverging) indicates that presence of a self-stabilization mechanism Damian et al. [2023] (which is absent in a quadratic) that regularizes the sharpness near $2/\rho$. The common occurrence of this phenomenon across various tasks, architectures and initialization has inspired substantial research on edge of stability. For example EoS has been studied across several non-convex optimization problems [Wang et al., 2022b, 2023a, Zhu et al., 2023, Chen and Bruna, 2023, Agarwala et al., 2023, Arora et al., 2022, Lyu et al., 2022, Ahn et al., 2024, Wu et al., 2024, Ghosh et al., 2025, Even et al., 2024, Chen et al., 2024, Kreisler et al., 2023, Kalra et al., 2025, Zhu et al., 2024] and across other descent based optimizers such as SGD [Lee and Jang, 2023, Andreyev and Beneventano, 2024], momentum [Phunyaphibarn et al., 2024], sharpness aware minimization (SAM) Agarwala and Dauphin [2023], Long and Bartlett [2024] and Adam/RMSProp [Cohen et al., 2022, 2025].
Stability of GD with Noise
Wu et al. [2018] analyze the linear dynamical stability of SGD, demonstrating that the batch size and the gradient covariance matrix impose an additional constraint, requiring the Hessian operator norm to be smaller than $2/\rho$. Wu et al. [2022] extend this work by assuming alignment between the gradient covariance matrix and the Fisher matrix, arguing that noise concentrates in the sharp directions of the landscape. Although neither study analyzes the EoS limit for SGD, they were instrumental in understanding the local stability of SGD near a global minimum. Lee and Jang [2023], Andreyev and Beneventano [2024] investigate the dynamics of SGD at EoS and instead propose that a different metric, called mini-batch aware sharpness, must be smaller than $2/\rho$ for stability. Since the actual sharpness is less than this mini-batch aware sharpness, sharpness itself must be smaller than $2/\rho$. However, in our work, the perturbation is in the weight-space and not on the gradient.
Figure 10: Shows the mechanism through which VL finds flatter minima. GD (gray iterates) get stuck in a local sharp minima whereas VL (red iterates) escapes to a flatter minima.
Appendix B Proof of Theorem 1
[…]
Proof.
We analyze the stability of the weight-perturbed gradient descent (GD) update applied to the quadratic loss
$\ell(\boldsymbol{m})=\frac{1}{2}\boldsymbol{m}^{\top}\mathbf{Q}\boldsymbol{m},$ (13)
where $\mathbf{Q}\in\mathbb{R}^{d\times d}$ is positive definite. The update rule is given by
$\boldsymbol{m}_{t+1}\leftarrow\boldsymbol{m}_{t}-\rho\,\hat{\boldsymbol{g}},\quad\text{where}\quad\hat{\boldsymbol{g}}:=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\nabla\ell(\boldsymbol{m}_{t}+\boldsymbol{\epsilon}_{i}),\quad\boldsymbol{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma}).$ (14)
Since $\ell$ is quadratic, its gradient at perturbed point satisfies
$\nabla\ell(\boldsymbol{m}_{t}+\boldsymbol{\epsilon})=\nabla\ell(\boldsymbol{m}_{t})+\mathbf{Q}\boldsymbol{\epsilon}.$ (15)
Therefore, if $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})$, then
$\nabla\ell(\boldsymbol{m}_{t}+\boldsymbol{\epsilon})\sim\mathcal{N}\left(\nabla\ell(\boldsymbol{m}_{t}),\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q}\right),$ (16)
by linear transformation of Gaussian variables.
As $\hat{\boldsymbol{g}}$ is the average of $N_{s}$ i.i.d. samples from this distribution, it follows that
$\hat{\boldsymbol{g}}\sim\mathcal{N}\left(\nabla\ell(\boldsymbol{m}_{t}),\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q}\right).$ (17)
Now, the Taylor expansion (exact) of the quadratic loss around $\boldsymbol{m}_{t}$ gives
$\ell(\boldsymbol{m}_{t+1})$ $=\ell(\boldsymbol{m}_{t})+\nabla\ell(\boldsymbol{m}_{t})^{\top}(\boldsymbol{m}_{t+1}-\boldsymbol{m}_{t})+\frac{1}{2}(\boldsymbol{m}_{t+1}-\boldsymbol{m}_{t})^{\top}\mathbf{Q}(\boldsymbol{m}_{t+1}-\boldsymbol{m}_{t}).$ (18)
Substituting the update rule (14) into (18), we get
$\ell(\boldsymbol{m}_{t+1})-\ell(\boldsymbol{m}_{t})$ $=-\rho\nabla\ell(\boldsymbol{m}_{t})^{\top}\hat{\boldsymbol{g}}+\frac{\rho^{2}}{2}\hat{\boldsymbol{g}}^{\top}\mathbf{Q}\hat{\boldsymbol{g}}$. (19)
We now observe that the change in loss, denoted by $\Delta\ell:=\ell(\boldsymbol{m}_{t+1})-\ell(\boldsymbol{m}_{t})$, is a random variable due to the stochasticity in the update. Using the expression from Equation (19), we can expand $\Delta\ell$ into a deterministic (mean) and a stochastic (fluctuation) component ($R$).
$\Delta\ell$ $=\ell(\mathbf{m}_{t+1})-\ell(\mathbf{m}_{t})=-\rho\,\mathbf{g}^{\top}\hat{\mathbf{g}}+\frac{\rho^{2}}{2}\hat{\mathbf{g}}^{\top}\mathbf{Q}\hat{\mathbf{g}}$ (20)
$=-\rho\,\mathbf{g}^{\top}(\mathbf{g}+\boldsymbol{\xi})+\frac{\rho^{2}}{2}(\mathbf{g}+\boldsymbol{\xi})^{\top}\mathbf{Q}(\mathbf{g}+\boldsymbol{\xi})$ (21)
$=\underbrace{-\rho\|\mathbf{g}\|^{2}+\frac{\rho^{2}}{2}\mathbf{g}^{\top}\mathbf{Q}\mathbf{g}+\frac{\rho^{2}}{2}\mathbb{E}[\boldsymbol{\xi}^{\top}\mathbf{Q}\boldsymbol{\xi}]}_{\mathbb{E}[\Delta\ell]}+\underbrace{\left(-\rho\,\mathbf{g}^{\top}\boldsymbol{\xi}+\rho^{2}\,\mathbf{g}^{\top}\mathbf{Q}\boldsymbol{\xi}+\frac{\rho^{2}}{2}\left(\boldsymbol{\xi}^{\top}\mathbf{Q}\boldsymbol{\xi}-\mathbb{E}[\boldsymbol{\xi}^{\top}\mathbf{Q}\boldsymbol{\xi}]\right)\right)}_{R}$ (22)
We write the perturbed gradient as $\hat{\mathbf{g}}=\mathbf{g}+\boldsymbol{\xi}$, where $\boldsymbol{\xi}\sim\mathcal{N}\left(\mathbf{0},\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q}\right)$. Here $\mathbf{g}$ denotes the gradient evaluated at the posterior mean, that is, $\nabla\ell(\boldsymbol{m}_{t})$. This expresses the update as the sum of a deterministic component $\mathbf{g}$ and a random fluctuation $\boldsymbol{\xi}$.
To ensure that the update leads to descent on average, we compute the expected change in loss and enforce the stability condition $\mathbb{E}[\Delta\ell]<0$ with respect to the curvature.
Taking the expectation of the loss change derived in Equation (19), we obtain:
$\mathbb{E}[\Delta\ell]=-\rho\mathbf{g}^{\top}(\mathbf{I}-\frac{\rho}{2}\mathbf{Q})\mathbf{g}+\frac{\rho^{2}}{2N_{s}}\operatorname{Tr}\left(\boldsymbol{\Sigma}\mathbf{Q}^{3}\right)<0$ (23)
Since the Hessian $\mathbf{Q}$ is symmetric, it admits an eigendecomposition $\mathbf{Q}=\sum_{i=1}^{d}\lambda_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\top}$, where $(\lambda_{i},\mathbf{v}_{i})$ are the eigenvalue-eigenvector pairs. We expand the expression in the eigenbasis of $\mathbf{Q}$. Note the following identities:
  • $\mathbf{g}^{\top}\mathbf{g}=\sum_{i=1}^{d}(\mathbf{v}_{i}^{\top}\mathbf{g})^{2}$,
  • $\mathbf{g}^{\top}\mathbf{Q}\mathbf{g}=\sum_{i=1}^{d}\lambda_{i}(\mathbf{v}_{i}^{\top}\mathbf{g})^{2}$,
  • $\operatorname{Tr}(\boldsymbol{\Sigma}\mathbf{Q}^{3})\leq\sum_{i=1}^{d}\sigma_{i}\lambda_{i}^{3}$  (Von Neuman’s trace inequality).
Thus, we obtain an upper bound on the expected descent:
$\mathbb{E}[\ell(\mathbf{m}_{t+1})]-\ell(\mathbf{m}_{t})$ $\leq\sum_{i=1}^{d}\left[-\rho(\mathbf{v}_{i}^{\top}\mathbf{g})^{2}+\frac{\rho^{2}}{2}\lambda_{i}(\mathbf{v}_{i}^{\top}\mathbf{g})^{2}+\frac{\rho^{2}}{2N_{s}}\sigma_{i}\lambda_{i}^{3}\right]$
$=:\sum_{i=1}^{d}f(\lambda_{i},\mathbf{v}_{i}).$
The inequality is an equality for an isotropic Gaussian $\boldsymbol{\Sigma}$. For anisotropic covariance matrix, the tightness of this inequality depends on the alignment of the Hessian and the covariance matrix.
To ensure descent in expectation, we require $f(\lambda_{i},\mathbf{v}_{i})<0$ for all $i$. We note that this is a sufficient condition for descent to take place. We further extend this Theorem to a necessary condition in Theorem. Define
$f(\lambda_{i})=-\rho a_{i}+\frac{\rho^{2}}{2}\lambda_{i}a_{i}+\frac{\rho^{2}}{2N_{s}}\sigma_{i}\lambda_{i}^{3},$
where $a_{i}:=(\mathbf{v}_{i}^{\top}\mathbf{g})^{2}>0$  [^1] For GD, if $\mathbf{v}_{i}^{T}\mathbf{m}_{0}=0$, $a_{i}=0$ always holds. However, stochasticity ensures that $a_{i}\neq 0$ even when an $\mathbf{m}_{0}$ is chosen such that it is orthogonal to some $\mathbf{v}_{i}$’s.. This is a cubic polynomial in $\lambda_{i}$ of the form
$f(\lambda)=a+b\lambda+c\lambda^{3},$
with:
$a$ $=-\rho a_{i}<0,$
$b$ $=\frac{\rho^{2}}{2}a_{i}>0,$
$c$ $=\frac{\rho^{2}}{2N_{s}}\sigma_{i}>0.$
Since $f^{\prime}(\lambda)=b+3c\lambda^{2}>0$ for all $\lambda>0$, the function is strictly increasing. Therefore, ensuring $f(\lambda_{i})<0$ is equivalent to requiring $\lambda_{i}$ to be smaller than the (unique) positive root of $f(\lambda)=0$.
By Cardano’s method for solving cubics, the condition $\Delta=\left(\frac{b}{3c}\right)^{3}+\left(\frac{a}{2c}\right)^{2}>0$ implies a unique real root. The root can be expressed as:
$\lambda_{i}$ $=\left(\frac{z_{i}}{\rho}\right)^{1/3}\left(\left(1+\sqrt{1+\frac{z_{i}\rho^{2}}{27}}\right)^{1/3}+\left(1-\sqrt{1+\frac{z_{i}\rho^{2}}{27}}\right)^{1/3}\right)$ (24)
$=2\sqrt{\frac{z_{i}}{3}}\sinh\left(\frac{1}{3}\sinh^{-1}\left(\frac{3}{\rho}\sqrt{\frac{3}{z_{i}}}\right)\right),$ (25)
where $z_{i}:=\frac{N_{s}a_{i}}{\sigma_{i}}=\frac{N_{s}(\mathbf{v}_{i}^{\top}\mathbf{g})^{2}}{\sigma_{i}}$.
Hence, the expected loss decreases if
$0<\lambda_{i}<2\sqrt{\frac{z_{i}}{3}}\sinh\left(\frac{1}{3}\sinh^{-1}\left(\frac{3}{\rho}\sqrt{\frac{3}{z_{i}}}\right)\right)\quad\text{for all }i.$
Now that we have established a sufficient condition on the curvature matrix $\mathbf{Q}$ to ensure that the expected loss decreases, we next address the random variability of the actual loss change due to the stochasticity of the update. Specifically, we show that the random variable $\Delta\ell$ concentrates sharply around its expectation.
For simplicity, we assumed to $\mathbf{Q}$ to be full rank. If $\mathbf{Q}$ has a non-trivial null-space that is $\lambda_{i}=0$ for some $d>i\geq k$, then we have $\mathbf{v}_{i}^{\top}\mathbf{g}=\mathbf{v}_{i}^{\top}\mathbf{Q}\mathbf{m}_{t}=\lambda_{i}\mathbf{v}_{i}^{\top}\mathbf{m}_{t}=0$. This means there is no component of the iterate in the null-space direction of $\mathbf{Q}$ and does not contribute to the descent objective.
This guarantees that, under the sufficient condition derived above, the actual loss decrease also holds with high probability, not just in expectation. The following lemma provides a concentration bound for $\Delta\ell$ about its mean:
[…]
Proof.
In Theorem 5, we showed that the descent step occurs in expectation, if the sharpness is less than the stability threshold. In this lemma, we show that the descent step occurs with high probability under the same condition. To show, this we use derive a concentration inequality to show that the descent step concentrates around its expectation with high probability. Moreover, with larger MC samples $N_{s}$, the deviation is small. The proof occurs in the following steps:
  1. 1.
    Let $\boldsymbol{\epsilon}=\hat{\mathbf{g}}-\mathbf{g}$ denote the gradient noise which is a rv $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q})$. The descent step has both linear and quadratic terms wrt $\boldsymbol{\epsilon}$.
  2. 2.
    To ensure the deviation of the descent $\Delta\ell$ from its expectation $\mathbb{E}[\Delta\ell]$, we analyze the concentration of both the linear term and the quadratic term. The concentration of the linear term follows simply from the Gaussian tail. The concentration of the quadratic term is done using the Hanson Wright concentration inequality Rudelson and Vershynin [2013], Vershynin [2018].
  3. 3.
    The tail bounds from both the linear and the quadratic term is combined using the union bound which concludes the proof.
Step-1: Separating fluctuation and expectation term
The descent step $\Delta\ell$ as derived in Theorem 5 can be written as its expectation $\mathbb{E}[\Delta\ell]$ and fluctuation $R$.
$\Delta\ell=l(\mathbf{m}_{t+1})-l(\mathbf{m}_{t})=-\rho\mathbf{g}^{T}\hat{\mathbf{g}}+\frac{\rho^{2}}{2}\hat{\mathbf{g}}^{T}\mathbf{Q}\hat{\mathbf{g}}$
$=-\rho\mathbf{g}^{T}(\mathbf{g}+\boldsymbol{\epsilon})+\frac{\rho^{2}}{2}(\mathbf{g}+\boldsymbol{\epsilon})^{T}\mathbf{Q}(\mathbf{g}+\boldsymbol{\epsilon})$
$=\underbrace{-\rho\|\mathbf{g}\|^{2}+\frac{\rho^{2}}{2}\mathbf{g}^{T}\mathbf{Q}\mathbf{g}+\frac{\rho^{2}}{2}\mathbb{E}[\boldsymbol{\epsilon}^{T}\mathbf{Q}\boldsymbol{\epsilon}}_{\mathbb{E}[\Delta\ell]}]+\underbrace{(-\rho\mathbf{g}^{T}\boldsymbol{\epsilon}+\rho^{2}\mathbf{g}^{T}\mathbf{Q}\boldsymbol{\epsilon}+\frac{\rho^{2}}{2}(\boldsymbol{\epsilon}^{T}\mathbf{Q}\boldsymbol{\epsilon}-\mathbb{E}[\boldsymbol{\epsilon}^{T}\mathbf{Q}\boldsymbol{\epsilon}]))}_{R}$
We show that the fluctuation $R=\Delta\ell-\mathbb{E}[\Delta\ell]$ is small with high probability, i.e, it has a sub-Gaussian tail.
Step-2: Concentration bound on the fluctuation
We separate the fluctuation on linear and quadratic terms wrt $\boldsymbol{\epsilon}$ since applying concentration inequality on each term is different.
$R=L+Q$, where $L=-\rho\mathbf{g}^{T}\boldsymbol{\epsilon}+\rho^{2}\mathbf{g}^{T}\mathbf{Q}\boldsymbol{\epsilon}$ and $Q=\frac{\rho^{2}}{2}(\boldsymbol{\epsilon}^{T}\mathbf{Q}\boldsymbol{\epsilon}-\mathbb{E}[\boldsymbol{\epsilon}^{T}\mathbf{Q}\boldsymbol{\epsilon}])$.
Tail bound on L using sub-Gaussin concentration: Since $L$ is a linear function of the Gaussian vector $\boldsymbol{\epsilon}$, it itself is Gaussian. its variance is
$\sigma^{2}_{L}=\frac{1}{N_{s}}\mathbf{h}^{T}(\mathbf{Q}\mathbf{\Sigma}\mathbf{Q})\mathbf{h}=\frac{\alpha_{L}}{N_{s}}$
where $\alpha_{L}:=\mathbf{h}^{T}\!\bigl(\mathbf{Q}\mathbf{\Sigma}\mathbf{Q}\bigr)\mathbf{h}$ and $\mathbf{h}=\rho^{2}\mathbf{Q}\mathbf{g}-\rho\mathbf{g}$. A standard Gaussian tail then implies that for any $t>0$,
$Pr\bigl(|L|\geq t\bigl)\leq 2\exp{\bigl(-\frac{t^{2}}{2\sigma^{2}_{L}}\bigl)}=2\exp{\bigl(-\frac{t^{2}N_{s}}{2\alpha_{L}}\bigl)}$
Choosing $t=\frac{\epsilon}{2}$ yields
$Pr\bigl(|L|\geq\frac{\epsilon}{2}\bigl)\leq 2\exp{\bigl(-\frac{\epsilon^{2}N_{s}}{8\alpha_{L}}\bigl)}$ (26)
Tail bound on Q using Hanson-Wright Concentration inequality: For the quadratic term, note that $Q=\frac{\rho^{2}}{2}(\boldsymbol{\epsilon}^{T}\mathbf{Q}\boldsymbol{\epsilon}-\mathbb{E}[\boldsymbol{\epsilon}^{T}\mathbf{Q}\boldsymbol{\epsilon}])$. The Hanson-Wright inequality states that a sub-Gaussian vector $\boldsymbol{\epsilon}$ (here Gaussian) with sub-Gaussian norm bounded by $K$ for any matrix $\mathbf{A}$ (in our case $\mathbf{A}=\frac{\rho^{2}}{2}\mathbf{Q}$), there exists universal constant $c_{1}>0$ such that for any $t>0$,
$\Pr\Bigl(\big|\boldsymbol{\epsilon}^{T}\mathbf{A}\boldsymbol{\epsilon}-\mathbb{E}[\boldsymbol{\epsilon}^{T}\mathbf{A}\boldsymbol{\epsilon}]\big|>t\Bigr)\leq 2\exp\Biggl(-c_{1}\min\Bigl\{\frac{t^{2}}{K^{4}\|\mathbf{A}\|_{F}^{2}},\frac{t}{K^{2}\|\mathbf{A}\|_{2}}\Bigr\}\Biggr).$
Since $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\frac{1}{N_{s}}\mathbf{Q}\mathbf{\Sigma}\mathbf{Q})$, its sub-Gaussian norm satisfies $K^{2}\leq c^{2}_{2}\lambda_{max}\bigl(\frac{1}{N_{s}}\mathbf{Q}\mathbf{\Sigma}\mathbf{Q}\bigl)=\frac{c^{2}_{2}\beta}{N_{s}}$, where we define $\beta=\lambda_{max}\bigl(\mathbf{Q}\mathbf{\Sigma}\mathbf{Q}\bigl)$ and $c_{2}$ is another universal constant. Furthermore, we have $K^{4}<\frac{c^{2}_{4}\beta^{2}}{N^{2}_{s}}$, $\|\mathbf{A}\|=\frac{\rho^{2}}{2}\|\mathbf{Q}\|$ and $\|\mathbf{A}\|_{F}=\frac{\rho^{2}}{2}\|\mathbf{Q}\|_{F}$. To substitute the sub-gaussian norm $K$ from the bound, we use the inequalities $\frac{t^{2}}{K^{4}\|\mathbf{A}\|^{2}_{F}}\geq\frac{t^{2}N^{2}_{s}}{c^{4}_{2}\beta^{2}\frac{\rho^{4}}{4}\|\mathbf{Q}\|^{2}_{F}}$ and $\frac{t}{K^{2}\|\mathbf{A}\|}\geq\frac{tN_{s}}{c^{2}_{2}\beta\frac{\rho^{2}}{2}\|\mathbf{Q}\|}$, we get the tail bound to be
$\Pr\Bigl(\big|Q|>t\Bigr)\leq 2\exp\Biggl(-c_{1}\min\Bigl\{\frac{t^{2}N^{2}_{s}}{c^{4}_{2}\beta^{2}\frac{\rho^{4}}{4}\|\mathbf{Q}\|^{2}_{F}},\frac{tN_{s}}{c^{2}_{2}\beta\frac{\rho^{2}}{2}\|\mathbf{Q}\|}\Bigr\}\Biggr).$
Finally chosing $c_{3}=\max\{c^{4}_{2}\beta^{2}\frac{\rho^{4}}{4}\|\mathbf{Q}\|^{2}_{F},c^{2}_{2}\beta\frac{\rho^{2}}{2}\|\mathbf{Q}\|\}$ and $t=\frac{\epsilon}{2}$, we get:
$\Pr\Bigl(\big|Q\big|>\frac{\epsilon}{2}\Bigr)\leq 2\exp\Biggl(-c_{1}\min\Bigl\{\frac{(\frac{\epsilon}{2})^{2}N^{2}_{s}}{c_{3}},\frac{(\frac{\epsilon}{2})N_{s}}{c_{3}}\Bigr\}\Biggr).$ (27)
Step-3: Combining $L$ and $Q$ concentation using union bound
Since $R=L+Q$, by the union bound and using equation (27) and (26) we get
$\Pr\Bigl(\big|R\big|>\epsilon\Bigr)\leq\Pr\Bigl(\big|L\big|>\frac{\epsilon}{2}\Bigr)+\Pr\Bigl(\big|Q\big|>\frac{\epsilon}{2}\Bigr)$
$\implies\Pr\Bigl(\big|R\big|>\epsilon\Bigr)\leq 2\exp{\bigl(-\frac{\epsilon^{2}N_{s}}{8\alpha_{L}}\bigl)}+2\exp\Biggl(-c_{1}\min\Bigl\{\frac{(\frac{\epsilon}{2})^{2}N^{2}_{s}}{c_{3}},\frac{(\frac{\epsilon}{2})N_{s}}{c_{3}}\Bigr\}\Biggr)$
We combine the sum to a single exponential tail by using the inequality $\exp(-X)+\exp(-Y)\leq 2\exp(-\min\{X,Y\})$. Now chosing $d_{2}=32\max\{\alpha_{L},c_{3}\}=32\max\{\alpha_{L},c^{4}_{2}\beta^{2}\frac{\rho^{4}}{4}\|\mathbf{Q}\|^{2}_{F},c^{2}_{2}\beta\frac{\rho^{2}}{2}\|\mathbf{Q}\|\}$ (substituting $c_{3}$) and $d_{1}=\min\{\frac{1}{32\alpha_{L}},\frac{c_{1}}{4}\}$ and using $\exp(-X)+\exp(-Y)\leq 2\exp(-\min\{X,Y\})$, we get
$\Pr\Bigl(\big|R\big|>\epsilon\Bigr)\leq 2\exp\Biggl(-d_{1}\min\Bigl\{\frac{\epsilon^{2}N^{2}_{s}}{d_{2}},\frac{\epsilon N_{s}}{d_{2}}\Bigr\}\Biggr)$
Assume the expected descent step is strictly negative by some margin $\delta>0$, i.e. $\mathbb{E}[\Delta\ell]\leq-\delta<0.$ Recall that $R=\Delta\ell-\mathbb{E}[\Delta\ell].$ Under this assumption, if $\Delta\ell\geq 0,$ then $\Delta\ell-\mathbb{E}[\Delta\ell]\geq\delta.$ Hence
$\{\Delta\ell\geq 0\}~\subseteq~\bigl\{\lvert R\rvert\geq\delta\bigr\}.$
We set $\epsilon=\delta$ in the concentration bound
$\Pr\!\bigl(\lvert R\rvert\geq\epsilon\bigr)~\leq~2\,\exp\!\Bigl(-\,d_{1}\,\min\Bigl\{\tfrac{\epsilon^{2}\,N_{s}^{2}}{d_{2}},\,\tfrac{\epsilon\,N_{s}}{d_{2}}\Bigr\}\Bigr),$
and obtain
$\Pr\!\bigl(\Delta\ell\geq 0\bigr)~\leq~2\,\exp\!\Bigl(-\,d_{1}\,\min\Bigl\{\tfrac{\delta^{2}\,N_{s}^{2}}{d_{2}},\,\tfrac{\delta\,N_{s}}{d_{2}}\Bigr\}\Bigr).$
Therefore, with probability at least
$1~-\;2\,\exp\!\Bigl(-\,d_{1}\,\min\Bigl\{\tfrac{\delta^{2}\,N_{s}^{2}}{d_{2}},\,\tfrac{\delta\,N_{s}}{d_{2}}\Bigr\}\Bigr),$
we have $\Delta\ell<0$. Thus, if the expected descent step is at most $\delta$, then $\Delta\ell$ is negative with high probability. ∎
Appendix C Role of posterior samples on a quadratic
On a quadratic loss, the perturbed averaged gradient follows a Gaussian distribution with variance inversely proportional to the number of posterior samples:
$\hat{\boldsymbol{g}}\sim\mathcal{N}\left(\nabla\ell(\boldsymbol{m}_{t}),\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q}\right).$ (28)
For large $N_{s}$, the dynamics of variational GD approach those of standard GD and exhibit similar stability characteristics. To examine the role of $N_{s}$, we consider a quadratic loss $\ell(m)=\frac{a}{2}m^{2}$ and compare GD with variational GD across varying values of $N_{s}$. When $\rho<2/a$, standard GD converges to the minimum. In the limit $N_{s}\to\infty$, variational GD recovers this behavior. However, for finite $N_{s}$, the gradient estimate becomes noisier, reducing the stability threshold as predicted by Theorem-1. Figure11 shows the resulting trajectories and histograms of the iterates. As $N_{s}$ decreases, the iterates exhibit greater variability and cover a wider range, reflecting increased instability. These results confirm the theoretical prediction that smaller $N_{s}$ increases the likelihood of the loss increasing in the next step, even under a stable learning rate.
(GD $\rho<\frac{2}{a}$)
$N_{s}=1000$
$N_{s}=200$
$N_{s}=100$
$N_{s}=50$
$N_{s}=20$
$N_{s}=10$
$N_{s}=5$
Figure 11: Histogram plot of iterates on a quadratic for GD with noise injection from a posterior distribution with $N_{s}$ samples. As $N_{s}$ decreases, the iterates become more unstable.
We further visualize the stability threshold predicted by Theorem-1 in Figure12. For each posterior sample size $N_{s}$, we compute the descent probability on a quadratic loss and plot it alongside the corresponding stability threshold $2/\rho\cdot\mathrm{VF}$. As shown in the top row, descent occurs with high probability whenever the curvature is below the threshold—consistent with Theorem 1. In the bottom row, we binarize the descent probability by setting it to 1 if it exceeds 0.5, and 0 otherwise. The resulting transition boundary closely aligns with the predicted threshold, further validating our theoretical result.
$N_{s}=1$
$N_{s}=2$
$N_{s}=5$
$N_{s}=10$
$N_{s}=50$
$N_{s}=1$
$N_{s}=2$
$N_{s}=5$
$N_{s}=10$
$N_{s}=50$
Figure 12: (Top Row): Probability of descent evaluated over 10 trials on several simulations of combination of quadratic curvature and noise variance. (Bottom Row): Hard thresholded probability for descent. The transition occurs near the boundary given by the derived expression of $2/\rho\cdot\text{VF}$ from Theorem-1.
Appendix D Loss Smoothing Beyond Local Quadratic Approximation
In our paper, we consider a local quadratic approximation of the loss to characterize the distribution of the perturbed gradient. By the property that a Gaussian distribution is invariant under linear transformation, we showed that the perturbed gradient (or it’s sample average) is also Gaussian. This is observed by characterizing the distribution of the perturbed gradient $\nabla\ell(\mathbf{m}_{t}+\boldsymbol{\epsilon})$. Employing a Taylor expansion around the current mean parameter $\mathbf{m}$, we obtain:
$\nabla\ell(\boldsymbol{m}+\boldsymbol{\epsilon})=\nabla\ell(\boldsymbol{m})+\nabla^{2}\ell(\boldsymbol{m})\boldsymbol{\epsilon}+\frac{1}{2}\nabla^{3}\ell(\boldsymbol{m})[\boldsymbol{\epsilon},\boldsymbol{\epsilon}]+O(\|\boldsymbol{\epsilon}\|^{3}),\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})$ (29)
In general, the distribution of $\nabla\ell(\boldsymbol{m}+\boldsymbol{\epsilon})$ involves higher-order terms of Gaussian variables, making exact characterization challenging. However, if the perturbation covariance $\boldsymbol{\Sigma}$ is sufficiently small relative to the local curvature, defined by the Hessian $\mathbf{H}(\boldsymbol{m})=\nabla^{2}\ell(\boldsymbol{m})$, specifically, when $\|\nabla^{3}\ell(\boldsymbol{m})\|_{op}\cdot\|\boldsymbol{\Sigma}\|_{2}\ll\|\mathbf{H}(\boldsymbol{m})\|_{2}$, the gradient approximation effectively simplifies to a linear approximation $\nabla\ell(\boldsymbol{m}+\boldsymbol{\epsilon})\approx\nabla\ell(\boldsymbol{m})+\mathbf{H}(\boldsymbol{m})\boldsymbol{\epsilon}$. This approximation is exact for a quadratic. Under this condition, since $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})$, the perturbed gradient is Gaussian $\nabla\ell(\boldsymbol{m}+\boldsymbol{\epsilon})\sim\mathcal{N}(\nabla\ell(\boldsymbol{m}),\mathbf{H}(\boldsymbol{m})\boldsymbol{\Sigma}\mathbf{H}(\boldsymbol{m})^{\top}).$
However, in high-perturbation regimes (large covariance $\Sigma$ where $\|\nabla^{3}\ell(\boldsymbol{m})\|_{op}\cdot\|\boldsymbol{\Sigma}\|_{2}\approx\|\mathbf{H}(\boldsymbol{m})\|_{2}$), the perturbed gradient is no more Gaussian (since it involves second-order terms on $\boldsymbol{\epsilon}$). Furthermore, the expectation of the perturbed gradient has a contribution from the neighborhood of the loss, which is referred to here as third-order curvature bias.
$\mathbb{E}[\nabla\ell(\mathbf{m}+\boldsymbol{\epsilon})]\approx\nabla\ell(\mathbf{m})+\tfrac{1}{2}\underbrace{\mathrm{Tr}_{2,3}(\boldsymbol{\Sigma}\nabla^{3}\ell(\mathbf{m}))}_{\text{third-order curvature bias}}+\mathcal{O}(\|\boldsymbol{\Sigma}\|^{2}).$ (30)
Having a large number of posterior samples, can help recover this modified expectation better with increasing number of samples. Here, the algorithm effectively follows the gradient of a smoothed version of the loss landscape, since samples are drawn from a wide neighborhood around $\boldsymbol{m}$. Accurately estimating this biased but meaningful descent direction under heavy-tailed noise requires a sufficiently large $N_{s}$, not just to reduce variance, but to faithfully recover the expected smoothed gradient.
In the next section and the subsequent theorem, we study this phenomenon in a non-quadratic function. Here we show that, first for a quadratic function, smoothing does not change the curvature of the underlying loss. But for a quartic function, smoothing by expectation changes the underlying curvature by the variance of the posterior, and furthermore, for smoothing by finite averaging, the curvature changes as a function of both the variance and the number of samples used to approximate the expectation.
Let $\ell(\theta)=a\theta^{2}+b\theta+c$ and let $q=\mathcal{N}(\theta|m,\sigma^{2})$ , then
$\ell_{conv}(m)=\mathbb{E}_{q}\ell(\theta)=\mathbb{E}_{q}(a\theta^{2}+b\theta+c)=am^{2}+bm+c+a\sigma^{2}$
So w.r.t the reparameterized variable $m$, the curvature of the loss remains unchanged, only shift occurs, proportional to $\sigma^{2}$. So, the stability dynamics on $\ell(\theta)$ and $\ell_{conv}(m)$ are the same. The loss diverges only when the learning rate is $\rho>\frac{2}{a}$. However, this is not the case with general losses, especially for losses where $\ell^{(4)}(\theta)\neq 0$, i.e if it has a non-zero fourth order derivative.
For example, let’s take the example of a quartic function $\ell(\theta)=(\theta^{2}-1)^{2}$ where minima is at $\theta^{*}=\pm 1$ and $\ell^{{}^{\prime\prime}}(\theta^{*})=12{\theta^{*}}^{2}-4=8$. Now, the smoothed loss $\ell_{conv}(m)=\mathbb{E}_{q}\ell(\theta)=\mathbb{E}_{q}(\theta^{2}-1)^{2}=m^{4}+m^{2}(6\sigma^{2}-2)+(3\sigma^{4}-2\sigma^{2}+1)$. The new loss has a global minima at $m^{*}=\pm\sqrt{1-3\sigma^{2}}$ (only if $\sigma<\sqrt{\frac{1}{3}}$). Comparing the curvature of both the losses at the global minima, we have:
$\ell^{{}^{\prime\prime}}(\theta)=12{\theta}^{2}-4=8$
$\ell_{conv}^{{}^{\prime\prime}}(m)=12{m}^{2}+2(6\sigma^{2}-2)=12m^{2}-4+12\sigma^{2}=8-24\sigma^{2}$
So, the smoothing operator, changes the curvature at the global minima for quartic functions. note that this wasn’t the case for quadratic functions, where the curvature was unchanged. This further changes the stability dynamics since the learning rate required for stable convergence is difference, for original loss $\ell(\theta)$, it is $\rho<\frac{2}{8}$, whereas for the smoothed loss, it is $\rho<\frac{2}{8-24\sigma^{2}}$. However, consider the effect of averaging of loss using finite mc samples $N_{s}$, $\ell_{avg}(\theta)=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\ell(\theta+\epsilon_{i})$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$:
$\ell_{avg}(\theta)=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\ell(\theta+\epsilon_{i})$
$=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}((\theta+\epsilon_{i})^{2}-1)^{2}$
$=\theta^{4}+\theta^{2}\left(\frac{1}{N_{s}}\sum_{i}^{N_{s}}(6\epsilon^{2}_{i}+4\epsilon_{i}-2)\right)+4\theta\left(\frac{1}{N_{s}}\sum_{i}^{N_{s}}\epsilon_{i}(\epsilon^{2}_{i}-1)\right)+\left(\frac{1}{N_{s}}\sum_{i}^{N_{s}}(\epsilon^{4}_{i}-2\epsilon^{2}_{i}+1)\right)$
So, the second derivative becomes:
$\ell^{{}^{\prime\prime}}_{avg}(\theta)=12\theta^{2}-4+\frac{2}{N_{s}}\sum_{i}^{N_{s}}(6\epsilon_{i}^{2}+4\epsilon_{i})$
If we compare the second derivative of the smoothed out-loss $\ell_{conv}^{{}^{\prime\prime}}(m)$ and averaged loss over $N_{s}$ samples $\ell^{{}^{\prime\prime}}_{avg}(\theta)$, we observe that wrt samples, the expectation is same, i.e,
$\mathbb{E}_{\epsilon_{i}}\ell^{{}^{\prime\prime}}_{avg}(\theta)=12\theta^{2}-4+\mathbb{E}_{\epsilon_{i}}[\frac{2}{N_{s}}\sum_{i}^{N_{s}}(6\epsilon_{i}^{2}+4\epsilon_{i})]=12\theta^{2}-4+12\sigma^{2}=\ell^{{}^{\prime\prime}}_{conv}(\theta)$
However, the pointwise deviation of $\ell^{{}^{\prime\prime}}_{avg}(\theta)$ and $\ell^{{}^{\prime\prime}}_{conv}(\theta)$ depends on the number of samples $N_{s}$ Taking the difference we have:
$\ell_{avg}^{{}^{\prime\prime}}(\theta)-\ell_{conv}^{{}^{\prime\prime}}(\theta)\;=\;\frac{2}{N_{s}}\sum_{i=1}^{N_{s}}\Bigl[\,(6\,\epsilon_{i}^{2}+4\,\epsilon_{i})\;-\;6\,\sigma^{2}\Bigr].$
Define
$Y_{i}\;=\;(6\,\epsilon_{i}^{2}+4\,\epsilon_{i})\;-\;6\,\sigma^{2}.$
Since $\mathbb{E}[\epsilon_{i}^{2}]=\sigma^{2}$ and $\mathbb{E}[\epsilon_{i}]=0$, each $Y_{i}$ has mean zero:
$\mathbb{E}[Y_{i}]\;=\;0.$
Hence,
$\ell_{avg}^{{}^{\prime\prime}}(\theta)\;-\;\ell_{conv}^{{}^{\prime\prime}}(\theta)\;=\;\frac{2}{N_{s}}\sum_{i=1}^{N_{s}}Y_{i}.$
Note that $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ is sub-Gaussian, and $\epsilon_{i}^{2}$ is sub-exponential. Thus $Y_{i}=6\,\epsilon_{i}^{2}+4\,\epsilon_{i}$ is a linear combination of a sub-exponential and a sub-Gaussian variable, which remains sub-exponential. Shifting by the constant  $6\,\sigma^{2}$ does not affect sub-exponential parameters. Therefore, each $Y_{i}$ is sub-exponential.
Let $(v,b)$ be sub-exponential parameters for $Y_{i}$. By standard Bernstein-type concentration for sub-exponential random variables, there exists a universal constant $c>0$ such that for all $t>0$:
$\Pr\!\Bigl(\,\Bigl|\tfrac{1}{N_{s}}\sum_{i=1}^{N_{s}}Y_{i}\Bigr|\,\geq\,t\Bigr)\;\;\leq\;\;2\,\exp\!\Bigl(\,-\,c\,N_{s}\,\min\Bigl\{\tfrac{t^{2}}{v},\,\tfrac{t}{b}\Bigr\}\Bigr).$
Since
$\ell_{avg}^{{}^{\prime\prime}}(\theta)-\ell_{conv}^{{}^{\prime\prime}}(\theta)\;=\;\frac{2}{N_{s}}\sum_{i=1}^{N_{s}}Y_{i},$
we have
$\Bigl|\ell_{avg}^{{}^{\prime\prime}}(\theta)-\ell_{conv}^{{}^{\prime\prime}}(\theta)\Bigr|\;=\;\frac{2}{N_{s}}\,\Bigl|\sum_{i=1}^{N_{s}}Y_{i}\Bigr|\;=\;2\,\Bigl|\tfrac{1}{N_{s}}\sum_{i=1}^{N_{s}}Y_{i}\Bigr|.$
Hence, for any $\delta>0$,
$\Pr\!\Bigl(\,\bigl|\ell_{avg}^{{}^{\prime\prime}}(\theta)-\ell_{conv}^{{}^{\prime\prime}}(\theta)\bigr|\;\geq\;\delta\Bigr)\;=\;\Pr\!\Bigl(\,\Bigl|\tfrac{1}{N_{s}}\sum_{i=1}^{N_{s}}Y_{i}\Bigr|\;\geq\;\tfrac{\delta}{2}\Bigr)\;\;\leq\;\;2\,\exp\!\Bigl(\,-\,c\,N_{s}\,\min\Bigl\{\tfrac{(\delta/2)^{2}}{v},\,\tfrac{\delta/2}{b}\Bigr\}\Bigr).$
Plugging in $v=c_{1}\sigma^{4}$ and $b=c_{2}\sigma^{2}$, which are the subexponential norms in terms of the variance, we get for some constant $c_{1}$ and $c_{2}$:
$\Pr\!\Bigl(\,\bigl|\ell_{avg}^{{}^{\prime\prime}}(\theta)-\ell_{conv}^{{}^{\prime\prime}}(\theta)\bigr|\;\geq\;\delta\Bigr)\;\leq\;2\,\exp\!\Bigl(\,-\,c\,N_{s}\,\min\Bigl\{\tfrac{\delta^{2}}{4c_{1}\sigma^{4}},\;\tfrac{\delta}{2c_{2}\sigma^{2}}\Bigr\}\Bigr).$
This inequality shows that the finite-sample second derivative $\ell_{avg}^{{}^{\prime\prime}}(\theta)$ concentrates around the infinite-sample second derivative $\ell_{conv}^{{}^{\prime\prime}}(\theta)$ at an exponential rate in $N_{s}$. While they are identical in expectation, the above result quantifies their pointwise deviation with high probability. We formalize this observation for general analytical function which has continous derivatives.
Theorem D.1 (Concentration of Smoothed Curvature).
Let $f:\mathbb{R}\to\mathbb{R}$ be six-times continuously differentiable analytical function, with all derivatives up to order 6 bounded. Suppose $\sigma^{2}\sup_{x}f^{(6)}(x)\;<\;\sup_{x}f^{(4)}(x)$ and for i.i.d. samples $\{\epsilon_{i}\}_{i=1}^{N_{s}}$ with $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, define
$f_{\mathrm{avg}}(x)\;=\;\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}f\bigl(x+\epsilon_{i}\bigr),\quad f_{\mathrm{conv}}(x)\;=\;\mathbb{E}_{z\sim\mathcal{N}(0,\sigma^{2})}\bigl[f(x+z)\bigr].$
Then there exist universal constants $c,\,c_{1},\,c_{2}>0$ such that for any $\delta>0$,
$\Pr\!\Bigl(\,\bigl|f_{\mathrm{avg}}^{\prime\prime}(x)\;-\;f_{\mathrm{conv}}^{\prime\prime}(x)\bigr|\,\geq\,\delta\Bigr)\;\leq\;2\,\exp\!\Bigl(\,-\,c\,N_{s}\,\min\Bigl\{\tfrac{\delta^{2}}{4\,c_{1}\,\sigma^{4}},\;\tfrac{\delta}{2\,c_{2}\,\sigma^{2}}\Bigr\}\Bigr).$
Hence, $\bigl|f_{\mathrm{avg}}^{\prime\prime}(x)-f_{\mathrm{conv}}^{\prime\prime}(x)\bigr|$ concentrates around zero at an exponential rate in $N_{s}$.
A general extension of this proof can be done for functions which has a finite fourth order derivative and small sixth order derivative such that $\sigma^{2}\sup_{x}f^{6}(x)<\sup_{x}f^{4}(x)$. Starting with a sufficiently smooth function $f\in C^{k}$, we can expand $f(x+z)$ in a Taylor series around $x$. Let $z\sim\mathcal{N}(0,\sigma^{2})$. Then:
$f(x+z)\;=\;f(x)\;+\;z\,f^{\prime}(x)\;+\;\frac{z^{2}}{2!}\,f^{\prime\prime}(x)\;+\;\frac{z^{3}}{3!}\,f^{(3)}(x)\;+\;\cdots$
Since $z$ is Gaussian with zero mean, all the odd moments $\mathbb{E}[z]$, $\mathbb{E}[z^{3}]$, etc., vanish. Thus, when we take the expectation,
$f_{\mathrm{conv}}(x)\;=\;\mathbb{E}_{z\sim\mathcal{N}(0,\sigma^{2})}\bigl[f(x+z)\bigr]\;=\;f(x)\;+\;\frac{\sigma^{2}}{2}\,f^{\prime\prime}(x)+\;+\;\frac{\sigma^{4}}{8}\,f^{(4)}(x)\;+\;\frac{\sigma^{6}}{48}\,f^{(6)}(x)\;+\;\cdots$
In the above, we use the fact that $\mathbb{E}[z^{2}]=\sigma^{2}$, $\mathbb{E}[z^{4}]=3\,\sigma^{4}$, $\mathbb{E}[z^{6}]=15\,\sigma^{6}$, etc., and only the even powers contribute. The curvature of the original function then becomes:
$f^{{}^{\prime\prime}}_{\mathrm{conv}}(x)=f^{\prime\prime}(x)+\frac{\sigma^{2}}{2}\,f^{4}(x)+\frac{\sigma^{4}}{8}\,f^{6}(x)+..$
Under the condition that $\sigma^{2}\sup_{x}f^{6}(x)<\sup_{x}f^{4}(x)$, we get
$f^{{}^{\prime\prime}}_{\mathrm{conv}}(x)\approx f^{\prime\prime}(x)+\frac{\sigma^{2}}{2}\,f^{4}(x)$ (31)
The smoothing effect does change the second order curvature of the loss. Furthermore, approximating This approximation makes the noise due to the $N_{s}$ samples be a subexponential and similar result holds.
Appendix E Discussion on stability and descent for Variational GD
On Necessity and Sufficiency of descent:
A further critical difference lies in the nature of the descent conditions for GD versus VGD. For GD on a quadratic loss, the standard descent lemma provides a condition on the maximum eigenvalue $\lambda_{max}<2/\rho$ that is both necessary and sufficient for ensuring stability and monotonic descent. A violation of this single condition guarantees divergence.
For VGD, the analysis is more nuanced. The true necessary and sufficient condition for the expected loss to decrease is that the sum of all components in its eigen-decomposition must be negative, as shown below:
$\mathbb{E}[\Delta\ell]=-\rho\mathbf{g}^{\top}(\mathbf{I}-\frac{\rho}{2}\mathbf{Q})\mathbf{g}+\frac{\rho^{2}}{2N_{s}}\operatorname{Tr}\left(\boldsymbol{\Sigma}\mathbf{Q}^{3}\right)=\sum_{i=1}^{d}f(\lambda_{i},\mathbf{v}_{i})<0$ (32)
However, analyzing this sum can be intractable. Instead, our work derives a practical sufficient condition by requiring each term in the sum to be negative independently, i.e., $f(\lambda_{i},\mathbf{v}_{i})<0$ for all i. This is the condition presented in Theorem 1, which yields the stability limit $\lambda_{i}<2/\rho\cdot\text{VF}(z_{i})$ for each eigenvalue.
This theoretical framework is validated by our extensive experiments on MLPs and ResNets across various learning-rate and variance. We consistently find that the leading Hessian eigenvalues, $\lambda_{i}$, hover around their respective stability thresholds, $2/\rho\cdot\text{VF}(z_{i})$. This alignment provides strong empirical support for our sufficient condition, showing it is an active constraint that accurately describes the behavior of VGD in practice.
Mulayoff and Michaeli [2024] also derives a stability condition (see their Theorem 5) which is a necessary and sufficient condition for the stability of SGD, but their setting is fundamentally different from ours. Firstly, in their setting, the assumption is that the batch-size is drawn uniformly at random from the finite set of all possible data samples. Our work, in contrast, does not model noise from data sampling but rather from perturbations to the model’s weights, which we assume follow an explicit, continuous distribution (e.g., Gaussian). This leads to fundamentally different noise structures. In our VGD framework, the resulting gradient estimator is shown to follow a normal distribution whose covariance, $\frac{1}{N_{s}}\mathbf{Q}\boldsymbol{\Sigma}\mathbf{Q}$, is shaped by the Hessian ($\mathbf{Q}$), meaning noise is amplified in directions of high curvature. In the SGD paper, the gradient noise has a discrete distribution with a covariance determined by the dataset’s intrinsic variance.
Appendix F Additional experiments across diverse datasets
While in the manuscript, we presented experiments on CIFAR-10, here we present several additional results to support our theorem and claims.
Gaussian Variational GD: In Figure 13, we compare GD and GD with weight perturbation trained across three different architectures on SVHN dataset Netzer et al. [2011]. We arrive at a similar conclusion that with Gaussian weight perturbation achieves smaller sharpness and better test accuracy than just GD. Similar trend is also observed for FashionMNIST dataset in Figure 15. In Figure 16, we show that in deep neural networks, sharpness also depends on the number of posterior samples, as smaller samples lead to smaller sharpness. In Figure 14, we plot the Variational factor along with the sharpness in ResNet and show that they match closely.
Sharpness comparison for Cross Entropy Loss: For cross-entropy (CE) loss, the sharpness dynamics differs from those observed with mean-squared error (MSE) loss. As training progresses and model predictions become increasingly confident, the term $p_{i}(1-p_{i})$ in the Hessian approaches zero, driving the curvature and hence sharpness to vanish. This causes both GD and VL (or IVON) to converge to regions with negligible sharpness, although their transient behaviors differ. VL’s sharper stability bound results in smaller peak sharpness compared to GD.
Figure 17 illustrates this effect for a fully connected neural network trained on CIFAR-10 with CE loss. The left panel shows the training loss, while the right panel presents the evolution of sharpness over training steps. Although both methods eventually exhibit vanishing curvature, IVON achieves a substantially lower peak sharpness than GD. This highlights the improved stability and flatter convergence landscape induced by variational learning even under the CE loss.
Weight-perturbation from heavy-tail distribution: In the manuscript, we presented results on perturbation from a heavy-tailed Student-t distribution for MLP trained on CIFAR-10. In Figure 18, we plot the training dynamics across ViT and ResNet-20 models. Here, we observe that heavier tails (smaller $\alpha$) leads to smaller sharpness and better test accuracy.
Experiments in Vision Transformers: In Attention-based architectures, such as Vision Transformers, it has been widely observed that sharpness for GD often goes above the stability threshold $2/\rho$. This phenomenon has been widely studied [Zhai et al., 2023] as attention entropy collapse that makes training unstable in Vision Transformers, with sharp spikes in both the training and test accuracy. To mitigate such unstable behaviour, weight perturbation can be used, due to its sharpness reducing effect and thereby stabilizing training. For example, in Figure 20, we perform an ablation study of training ViTs with different perturbation covariance. Noise with larger variance consistently leads to more stable training and smaller sharpness in ViT. Similarly, for VON we observe that preconditioned sharpness is smaller than $2/\rho$ for ViT training 19.
Experiments in NLP tasks: To verify that Variational GD with isotropic Gaussian noise achieves lower sharpness compared to GD on an NLP task, we add a classification head to a frozen BERT-mini backbone and finetune the head on SST-2. For both GD and Variational GD, we use the same learning rate of 0.05. We report train loss, validation loss, and sharpness in Figure 21 and observe lower sharpness for Variational GD throughout. ∎
(a) ViT
(b) MLP
(c) ResNet-20
Figure 13: Similar to the trend in Fig. 4, weight perturbed GD consistently finds flatter minima on the SVHN dataset.
(a) $\rho=0.01$
(b) $\rho=0.02$
(c) $\rho=0.05$
Figure 14: Normalized Sharpness $\|\nabla^{2}\ell(\mathbf{m}_{t})\|_{2}/(2/\rho)$ hovers about the Variational Factor in ResNet.
(a) ViT
(b) MLP
(c) ResNet-20
Figure 15: Similar to the trend in Fig. 4, variational learning finds flatter minima on the FashionMNIST dataset.
Figure 16: Using more samples per iteration leads to a large sharpness, as demonstrated in our ablation study where an MLP network is trained on a subset of the CIFAR-10 dataset containing 10000 images.
Figure 17: Training dynamics of fully connected neural network trained with GD and IVON with fixed covariance. Although the sharpness always drops to zero fro CE loss, IVON achives smaller peak sharpness than GD.
(a) ResNet-20
(b) ViT
Figure 18: We notice similar trends shown in Fig. 7 in training ViT and ResNet models. That is to say, smaller $\alpha$ corresponds to heavier-tailed posterior which leads to smaller sharpness. Note that as $\alpha$ approaches infinity, the Gaussian posterior is recovered.
Figure 19: We notice similar trends shown in Fig. 8 in training ViT models. That is to say, Smaller temperatures which shrinks the posterior reaches larger preconditioned sharpness.
Figure 20: For ViT, GD training is prone to loss spikes, and the sharpness often goes above the stability threshold $2/\rho$. On the other hand, with weight perturbation the training becomes more stable, and larger noise variance consistently leads to smaller sharpness.
Figure 21: SST-2 head finetuning on a frozen BERT mini backbone comparing Variational GD with isotropic Gaussian noise to vanilla GD, both with learning rate 0.05.
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\lambda_{4}$
(e) $\lambda_{5}$
(f) $\lambda_{6}$
(g) $\lambda_{7}$
(h) $\lambda_{8}$
(i) $\lambda_{9}$
(j) $\lambda_{10}$
Figure 22: Eigenspectrum ($\lambda_{1}$-$\lambda_{10}$) vs corresponding stability threshold $2/\rho\cdot\text{VF}(z_{i})$. 2-layer MLP trained with VGD $\rho=0.05$ and $\sigma^{2}=0.1$.
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\lambda_{4}$
(e) $\lambda_{5}$
(f) $\lambda_{6}$
(g) $\lambda_{7}$
(h) $\lambda_{8}$
(i) $\lambda_{9}$
(j) $\lambda_{10}$
Figure 23: Eigenspectrum ($\lambda_{1}$-$\lambda_{10}$) vs corresponding stability threshold $2/\rho\cdot\text{VF}(z_{i})$. 2-layer MLP trained with VGD $\rho=0.02$ and $\sigma^{2}=0.5$.
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\lambda_{4}$
(e) $\lambda_{5}$
(f) $\lambda_{6}$
(g) $\lambda_{7}$
(h) $\lambda_{8}$
(i) $\lambda_{9}$
(j) $\lambda_{10}$
Figure 24: Eigenspectrum ($\lambda_{1}$-$\lambda_{10}$) vs corresponding stability threshold $2/\rho\cdot\text{VF}(z_{i})$. 2-layer MLP trained with VGD $\rho=0.02$ and $\sigma^{2}=1.0$.
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\lambda_{4}$
(e) $\lambda_{5}$
(f) $\lambda_{6}$
(g) $\lambda_{7}$
(h) $\lambda_{8}$
(i) $\lambda_{9}$
(j) $\lambda_{10}$
Figure 25: Eigenspectrum ($\lambda_{1}$-$\lambda_{10}$) vs corresponding stability threshold $2/\rho\cdot\text{VF}(z_{i})$. 2-layer MLP trained with VGD $\rho=0.1$ and $\sigma^{2}=1.0$.
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\lambda_{4}$
(e) $\lambda_{5}$
(f) $\lambda_{6}$
(g) $\lambda_{7}$
(h) $\lambda_{8}$
(i) $\lambda_{9}$
(j) $\lambda_{10}$
Figure 26: Eigenspectrum ($\lambda_{1}$-$\lambda_{10}$) vs corresponding stability threshold $2/\rho\cdot\text{VF}(z_{i})$. ResNet-20 trained with VGD $\rho=0.1$ and $\sigma^{2}=0.1$.
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\lambda_{4}$
(e) $\lambda_{5}$
(f) $\lambda_{6}$
(g) $\lambda_{7}$
(h) $\lambda_{8}$
(i) $\lambda_{9}$
(j) $\lambda_{10}$
Figure 27: Eigenspectrum ($\lambda_{1}$-$\lambda_{10}$) vs corresponding stability threshold $2/\rho\cdot\text{VF}(z_{i})$. ResNet-20 trained with VGD $\rho=0.1$ and $\sigma^{2}=0.5$ (low variance).
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\lambda_{4}$
(e) $\lambda_{5}$
(f) $\lambda_{6}$
(g) $\lambda_{7}$
(h) $\lambda_{8}$
(i) $\lambda_{9}$
(j) $\lambda_{10}$
Figure 28: Eigenspectrum ($\lambda_{1}$-$\lambda_{10}$) vs corresponding stability threshold $2/\rho\cdot\text{VF}(z_{i})$. ResNet-20 trained with VGD $\rho=0.05$ and $\sigma^{2}=0.1$.
(a) $\lambda_{1}$
(b) $\lambda_{2}$
(c) $\lambda_{3}$
(d) $\lambda_{4}$
(e) $\lambda_{5}$
(f) $\lambda_{6}$
(g) $\lambda_{7}$
(h) $\lambda_{8}$
(i) $\lambda_{9}$
(j) $\lambda_{10}$
Figure 29: Eigenspectrum ($\lambda_{1}$-$\lambda_{10}$) vs corresponding stability threshold $2/\rho\cdot\text{VF}(z_{i})$. ResNet-20 trained with VGD $\rho=0.05$ and $\sigma^{2}=0.5$.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Thorough experiments show that the observations made in the theorem perfectly show up in experiments.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We state that our work is only limited to weight perturbed GD but not preconditioning.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Theorem-1, which states the main result of the paper has been validated extensively through experiments. For example in Figure-2, the theoritically calculated thereshold matches the experiment.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The code and the details of reproduction such as hyperparameter setting has been provided.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide anonymous code for reviewers to check. Once paper is published we will release the original code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, all these details have been stated explicitly.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Yes, averaged over various random seeds.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, it states the computer resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, it does.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Impact section added.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification:
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, all code sources cited and credited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We added a Readme file to document our code so reviewers can check them.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification:
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [No]
  74. Justification:
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Not used. only grammar checks and writing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.